OpenStudy (anonymous):

I keep getting a math error every time I plug e into my calculator? Anyone know what I'm doing wrong?

Directrix (directrix):

What expression are you plugging in?

OpenStudy (anonymous):

I think it's the natural log

OpenStudy (mertsj):

If you want to get e, enter 1 and then push the e^x button

OpenStudy (anonymous):

I'm doing that, but i get a math error when I want to get the answer

OpenStudy (anonymous):

or e^x then 1, depending on the calculator

OpenStudy (mertsj):

Yes, good point. Try that.

OpenStudy (accessdenied):

you should provide calculator you're using + exact error msg and what you entered

OpenStudy (unklerhaukus):

~2.718281828459045

OpenStudy (anonymous):

Wait a bit.

OpenStudy (anonymous):

if your calculator opens a left parentheses, make sure to close that parentheses

OpenStudy (anonymous):

There's that many calculators out there it's kind of difficult to help with. I'd suggest searching Google for your calculator's manual, and trying to determine the appropriate method.

OpenStudy (anonymous):

I'm using Casio, I got a math error, and I plugged in 2500e^(0.035X8)

OpenStudy (unklerhaukus):

i bet there are two varieties of e on you calculator , you might need to first enter [ shift ] or something like that

OpenStudy (anonymous):

have you tried 2500 * e^(0.035*8)?

OpenStudy (anonymous):

Um, not yet. wait.

OpenStudy (anonymous):

This time it said syntax error... :/

OpenStudy (mertsj):

Try .035 * 8= Then e^x= Then*2500=

OpenStudy (anonymous):

Oh!!! I just got it, turns out I was pressing the shift key instead of the alpha key. Thanks to everyone anyway!

OpenStudy (anonymous):

No, 3307.82

OpenStudy (mertsj):

Yep. Good for you.